A solid metal cube of side 10cm is heated from 10°c to 60°c. If the linear expansivity of the metal is 1•2×10-⁵k-¹. Calculate the increase in its volume.


1
Expert's answer
2021-02-19T10:30:17-0500

We can find the increase in the volume of the cube from the formula:


"\dfrac{\Delta V}{V_0}=\alpha_V\Delta T,""\Delta V=3\alpha_LV_0\Delta T,""\Delta V=3\cdot1.2\cdot10^{-5}\ ^{\circ}C^{-1}\cdot (10\ cm)^3\cdot50\ ^{\circ}C=1.8\ cm^3."